# The reliability of hybrid functionals for accurate fundamental and optical gap prediction of bulk solids and surfaces

### v1

@@ Abstract
We repeat this set of calculations with a similar slab representing the reconstructed Ge(111)-(2$\times$1) surface (see right side of table LABEL:table1). We find the same general trends as in the Si case, but with a few caveats: First, as shown above, the fundamental gap of bulk Ge at $\Gamma$ is known to be almost zero for most semilocal and meta-GGA functionals. Second, even when using a hybrid, specifically HSE, as a starting point for the $GW$ calculations, an overestimation of the bulk Ge gap is seen. This shows the sometimes inconsistent nature of ‘one-shot’ $GW$, even with a hybrid functional starting point used in previous studies Rinke et al. (2005); Gant et al. (2022); Golze et al. (2019).
@@
Finally, using linear response TDDFT  Casida (1996); Tretiak and Chernyak (2003), within the Tamm-Dancoff approximation, it is possible to obtain the linear absorption spectra of the reconstructed surfaces, here in arbitrary units for the oscillator strength. The results of our calculations are seen in Fig. 1 for the reconstructed Si (c) and Ge (d) slabs for PBE0, HSE, and SRSH functionals. The attainment of the optical gap for these slabs, made clear by Fig. 1, goes beyond the expectations of Ref. Jain et al. (2011) from hybrid functionals, owing to the use of TDDFT rather than ground-state DFT. We find that both the SRSH and HSE functionals perform with good accuracy, and get within 0.08 and and 0.14 eV of the experimental optical gap for Si respectively (for Ge the errors are 0.13 eV and 0.16 eV, respectively). When compared to previously reported $GW$-BSE optical gaps for the Si case, the SRSH and HSE are within 0.12 and 0.1 eV (for Ge those errors are 0.15 and 0.14 eV) respectively. PBE0 tends to overestimate the optical gap for the Si and Ge surface. While the calculated spectra in Fig. 1 allow for a visual comparison of the TDDFT excited states associated with the surface states for different functionals, the magnitude of the peak depends on the supercell and slab size, and therefore are taken as arbitrary. The lowest-energy optical peaks for the Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) slabs, at the WOT-SRSH level, are well within the $\pm$ 0.2 eV experimental uncertainty and prior GW-BSE calculations Jain et al. (2011). In the SISup , we report the individual excitons that make up the absorbance peaks and analyze the effects of Lorenzian broadening. Other studies have shown that in-gap states induce novel low-lying excitons not present in the bulk, e.g., those originating with a mixture of bulk states and surface states Hernangómez-Pérez et al. (2023). A detailed study of surface excitons would be an interesting subject for future work.
@@
To conclude, we find that it is possible to reliably predict both the bulk and surface state fundamental gaps, of Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) with the SRSH (using previously tuned bulk $\alpha$, and $\gamma$ parameters) and HSE functionals, with good accuracy. More importantly, we find that it is possible to obtain accurate optical gaps, using these hybrid functionals, with TDDFT calculations. While both SRSH and HSE functionals perform equally well here, we expect to see deviations in the performance of larger gap materials, similar to the bulk case (e.g. Fig. 3 in SI of Wing et al. (2021), and Chen et al. (2018)). Previous work more than a decade ago Jain et al. (2011) concluded that hybrid functionals (at the time) did not give reliable band gaps for bulk systems or surfaces, and that fortuitously obtaining a good surface state and bulk gap with a hybrid functional did not guarantee the optical gap with ground state DFT calculation. Here, we overcome this by using the DFT and TDDFT frameworks with a new generation of SRSH, namely WOT-SRSH, where $\alpha$ and $\gamma$ are constrained by an IP ansatz, can lead to bulk and surface fundamental and optical gaps in excellent agreement with experiments and previous $GW$-BSE results. Finally, while a tuning procedure is not yet available for surfaces, the reliable performance seen with the transferability of bulk parameters for these reconstructed surfaces, provides an indication that the construction of this class of hybrids goes in the correct direction.
I Acknowledgements
@@
F.S. acknowledges discussions with Prof. M. Jain which clarified important aspects of past work. M.C.-G. and F.S. thank Prof. A. Ramasubramaniam for fruitful discussions. M.C.-G. is grateful to the Azrieli foundation for the award of an International Postdoctoral Fellowship. L.K. acknowledges additional support from the Aryeh and Mintzi Katzman Professorial Chair and from the Helen and Martin Kimmel Award for Innovative Investigation. The formalism used in this work was funded through NSF–Binational Science Foundation Grant No. DMR-2015991 and the Israel Science Foundation; and our calculations were supported by the Theory of Materials Program at LBNL, funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Contract No. DE-AC02-05CH11231. Computational resources were provided by the NSF funded ACCESS program (frontera supercomputer),at the Texas Advanced Computing Center (TACC), the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility at Lawrence Berkeley National Laboratory, and the Chemfarm facility at the Weizmann Institute of Science.

### v2

@@ Abstract
We repeat this set of calculations with a similar slab representing the reconstructed Ge(111)-(2$\times$1) surface (see right side of table LABEL:table1). We find the same general trends as in the Si case, but with a few caveats: First, as shown above, the fundamental gap of bulk Ge at $\Gamma$ is known to be almost zero for most semilocal and meta-GGA functionals. Second, even when using a hybrid, specifically HSE, as a starting point for the $GW$ calculations, an overestimation of the bulk Ge gap is seen (yields a gap of $1.18$ eV). This shows the sometimes inconsistent nature of ‘one-shot’ $GW$, even with a hybrid functional starting point used in previous studies Rinke et al. (2005); Gant et al. (2022); Golze et al. (2019).
@@
Finally, using linear response TDDFT  Casida (1996); Tretiak and Chernyak (2003), within the Tamm-Dancoff approximation (TDA), it is possible to obtain the linear absorption spectra of the reconstructed surfaces, here in arbitrary units for the oscillator strength. The TDA has been found to be accurate for systems of similar excitonic effect, compared to the full Casida equation Sander et al. (2015); Byun and Ullrich (2017). The results of our calculations are seen in Fig. 1 for the reconstructed Si (c) and Ge (d) slabs for PBE0, HSE, and SRSH functionals. The attainment of the optical gap for these slabs, made clear by Fig. 1, goes beyond the expectations of Ref. Jain et al. (2011) from hybrid functionals, owing to the use of TDDFT rather than ground-state DFT. We find that both the SRSH and HSE functionals perform with good accuracy, and get within 0.08 and and 0.13 eV of the experimental optical gap for Si respectively (for Ge the errors are 0.13 eV and 0.16 eV, respectively). When compared to previously reported $GW$-BSE optical gaps for the Si case, the SRSH and HSE are within 0.12 and 0.17 eV (for Ge those errors are 0.15 and 0.14 eV) respectively. PBE0 tends to overestimate the optical gap for the Si and Ge surface. While the calculated spectra in Fig. 1 allow for a visual comparison of the TDDFT excited states associated with the surface states for different functionals, the magnitude of the peak depends on the supercell and slab size, and therefore are taken as arbitrary. The lowest-energy optical peaks for the Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) slabs, at the WOT-SRSH level, are well within the $\pm$ 0.2 eV experimental uncertainty and prior GW-BSE calculations Jain et al. (2011). In the SISup , we report the individual excitons that make up the absorbance peaks and analyze the effects of Lorentzian broadening. Other studies have shown that in-gap states induce novel low-lying excitons not present in the bulk, e.g., those originating with a mixture of bulk states and surface states Hernangómez-Pérez et al. (2023). A detailed study of surface excitons would be an interesting subject for future work.
@@
To conclude, we find that it is possible to reliably predict both the bulk and surface state fundamental gaps, of Si(111)-(2$\times$1) and Ge(111)-(2$\times$1) with the SRSH (using previously tuned bulk $\alpha$, and $\gamma$ parameters) and HSE functionals, with good accuracy. More importantly, we find that it is possible to obtain accurate optical gaps, using these hybrid functionals, with TDDFT calculations. While both SRSH and HSE functionals perform equally well here, we expect to see deviations in the performance of larger gap materials, similar to the bulk case e.g. Fig. 3 in SI of Wing et al. (2021), and Ref. Chen et al. (2018). Such comparisons involving HSE and PBE0 of bulk materials have also been studied e.g. Song et al. (2013), and table SIII.1 of Gant et al. (2022). Previous work more than a decade ago Jain et al. (2011) concluded that hybrid functionals (at the time) did not give reliable band gaps for bulk systems or surfaces, and that fortuitously obtaining a good surface state and bulk gap with a hybrid functional did not guarantee the optical gap with ground state DFT calculation. Here, we overcome this by using the DFT and TDDFT frameworks with a new generation of SRSH, namely WOT-SRSH, where $\alpha$ and $\gamma$ are constrained by an IP ansatz, can lead to bulk and surface fundamental and optical gaps in excellent agreement with experiments and previous $GW$-BSE results. Finally, while a tuning procedure is not yet available for surfaces, the reliable performance seen with the transferability of bulk parameters for these reconstructed surfaces, provides an indication that the construction of this class of hybrids goes in the correct direction.
I Supporting Information
Computational details, convergence tests, projected band structures for functionals not included in main text, and atomic structure details.
II Acknowledgements
@@
F.S. acknowledges discussions with Prof. M. Jain which clarified important aspects of past work. M.C.-G. and F.S. thank Prof. A. Ramasubramaniam for fruitful discussions. M.C.-G. is grateful to the Azrieli foundation for the award of an International Postdoctoral Fellowship and acknowledges support from the Diputación Foral de Gipuzkoa through Grant 2024-FELL-000007-01. F.R. acknowledges support from the BEWARE scheme of the Wallonia-Brussels Federation for funding under the European Commission’s Marie Curie-Skłodowska Action (COFUND 847587). L.K. acknowledges additional support from the Aryeh and Mintzi Katzman Professorial Chair and from the Helen and Martin Kimmel Award for Innovative Investigation. The formalism used in this work was funded through NSF–Binational Science Foundation Grant No. DMR-2015991 and the Israel Science Foundation; and our calculations were supported by the Theory of Materials Program at LBNL, funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Contract No. DE-AC02-05CH11231. Computational resources were provided by the NSF funded ACCESS program (frontera supercomputer),at the Texas Advanced Computing Center (TACC), the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility at Lawrence Berkeley National Laboratory, and the Chemfarm facility at the Weizmann Institute of Science.